# Magnetic field generation in mergers of massive main sequence stars

### v1

@@ 3.5 Dynamical ejecta
Because the further evolution of the merger product and also its optical display might be influenced by dynamical ejecta, we compute how much mass is unbound at the end of the simulation. To this end, we sum the mass of all cells in the simulation for which the total energy is positive. To get an upper limit, contributions from the potential, kinetic and also internal energies are taken into account. After the merger settled into the core–disk structure, at d, about of material are unbound, corresponding to 0.14% of the total mass of the system. Not much mass is ejected compared to other stellar mergers (e.g., common envelope phases, see Ohlmann et al. 2016a), also because the binding energy of the system is still quite large with about . The unbound mass moves outwards with an average radial velocity of about 180 km/s and is predominantly ejected near the orbital plane with a distribution that shows some tails to the polar directions.
@@ 4.1 Impact of numerical parameters
The main numerical parameters that affect the simulation outcome are the resolution, the initial distance and the magnetic field configuration. To assess the impact of the resolution, we ran model 1 with about 10 times more cells than model 2 (cf. Table 2). Because the adaptive mesh leads to cells with nearly constant mass, this leads to a higher resolution in volume by a factor of 10, which roughly corresponds to a factor of 2 in each spatial dimension. Overall, these two simulations show quite similar behavior in their orbital evolution (Fig. 2). The evolution of the magnetic field amplification differs slightly for the two models (Fig. 6): the amplification sets in later for model 2 with less cells. The final field strength and configuration, however, are quite similar which demonstrates that they are robust with respect to the resolution.
The initial distance influences the outcome as more angular momentum is present in the system for larger initial distances. In principle, the simulation should start at Roche-lobe overflow; however, this is prohibitively expensive in computational time because the mass transfer progresses only slowly in the beginning, leading to a long time to the actual merger. As already described, we remove angular momentum artificially in order to obtain a starting configuration that leads to a merger more quickly. In reality, the system would have more angular momentum, leading probably to a longer mass loss via L2 and L3 in the phase prior to the actual merger. The largest impact of this is probably on the detailed structure of the outer part of the disk. To assess the impact of a larger initial separation, we ran model 3 with a larger initial distance. As one can see in Fig. 2, it takes longer until the cores finally merge, but the overall evolution is similar. Also, the magnetic field evolution (Fig. 6) shows a slightly different shape, but the final field strength and configuration are similar. Thus, we expect that the main conclusions are not affected by the detailed choice of this parameter, but that it mainly affects the outermost structure of the disk.
@@
A typical numerical issue in MHD simulations is fulfilling the divergence constraint . In principle, the continuous equations of MHD preserve this constraint for initial conditions fulfilling it. However, discretization errors also lead to errors in . As explained in Pakmor & Springel (2013), we use the Powell method (Powell et al. 1999) to control the errors in the divergence. We have checked that the errors in our simulations behave very similarly as in the tests shown there (see Fig. 11): averages over larger regions are small and typically adjacent cells show different signs in the error, indicating that the large-scale impact of these local errors is small. The volume-weighted average of the relative divergence error is , , , and for the times given in Fig. 11.
@@ 4.2 Comparison to other simulations
Glebbeek et al. (2013) have conducted a series of stellar merger simulations as head-on collisions and subsequent stellar evolution calculations for a range of masses and evolutionary points. They used an SPH code with particles per star whereas we use a grid-based approach with 1 to 2 million cells per star. Moreover, we include MHD effects, which is the major novelty in our simulations. By simulating head-on collisions, they avoid dealing with rotation in the merger product and also with angular momentum transport. The collision may lead to more heating compared to the merger in our simulations. Moreover, there is probably no disk forming around the merger product in their simulations due to the missing angular momentum. In their simulations, the equal-mass collisions eject between 4% and 8% of the total mass (see their Figure 3), whereas in our simulation, only 0.14% of the mass is unbound (compare Sect. 3.5). This might be due to the difference in the initial conditions as material can escape more easily in head-on collisions, whereas in our merger simulations a large disk forms that is bound. In our simulation, the core is dominated by material from the secondary star, corresponding to “case S” in Glebbeek et al. (2013). As explained there, this can be attributed to the lower buoyancy in the secondary star. They find “case S” for relatively unevolved primaries at moderate mass ratios, which would also apply to our case.
@@
Glebbeek & Pols (2008) have analyzed stellar collision products. However, they do not conduct hydrodynamic simulations, but they use the entropy sorting method from Lombardi et al. (2002) to compute the merger products and then they focus on the further evolution. Lombardi et al. (2002) use results from earlier collision simulations using an SPH code to construct a model for predicting the outcome of such a collision. This model mainly sorts the material by its entropy but takes into account more effects such as mass loss and shock heating. For our simulations, comparing the merger product with an entropy-sorted mix of the initial stellar models, one can clearly see similar effects: the entropy in the simulated merger is larger especially in the center and in the outermost parts. Simple entropy sorting of the initial stellar models does not provide the right structure of the merged stars. Shock heating, chemical mixing and changes in the total mass of the merged star and thereby in density and temperature modify the entropy of the original fluid elements. Still, the final merged star has an outward increasing entropy profile as is expected for buoyantly stable stratifications (Landau & Lifshitz 1987). The models from Lombardi et al. (2002) are based on earlier collision simulations, amongst others the off-axis collisions conducted by Sills et al. (2001). They use an SPH code with about particles for the full simulation and target low-mass stars, mainly to study blue stragglers similar to Lombardi et al. (2002) and Glebbeek & Pols (2008).
@@
Ryu et al. (2024) simulate parabolic head-on collisions of low-mass stars with non-zero impact parameters similar to those mentioned above but they also use the moving-mesh MHD code Arepo and find strong magnetic field amplification. Such collisions contain less angular momentum than mergers in circular binaries studied here and thus form not so massive disk-like structures in the merger remnant as found here. The magnetic field amplification mechanism is otherwise in qualitative agreement with that found in our inspiral mergers of circular high-mass binaries: the collision mergers invoke several contact/grazing encounters before the final coalescence and in each phase small-scale dynamos amplify the magnetic fields. From the work presented in Ryu et al. (2024), however, it remains unclear whether also large-scale magnetic fields are formed. Vynatheya et al. (2025) compare head-on collisions of massive stars with binary mergers of the same stars. They find systematic differences between those, in particular they find that merger remnants have signs of an ordering of the magnetic field similar to our simulation, that is absent for collision remnants. They focus on a qualitative analysis of the final state of their collision and merger remnants. The latter seem consistent with our simulation.
@@
It is remarkable that the magnetic field amplification process in mergers of MS stars is so similar to that in mergers of white dwarfs and neutron stars presented in Pakmor et al. (2024) and Kiuchi et al. (2024), respectively. Also in these binary white-dwarf and neutron-star mergers small scale magnetic fields are produced by a turbulent dynamo and magnetic energy is then transferred to larger scales by a large-scale, MRI-driven dynamo. A similar situation is also found in neutron star–white dwarf mergers (Morán-Fraile et al. 2024) and common-envelope phases of low-mass and high-mass stars (Ondratschek et al. 2022; Vetter et al. 2024). Despite the similarities, all these simulations feature outflows that are magnetically driven, whereas our simulations do not show such outflows. This might be because we could not follow our simulations for long enough after the merger.
@@ 5 Conclusions
We have studied in detail the magnetic-field amplification in the merger of a and main-sequence star in an isolated binary system and the properties of the resulting merged star. The simulations were first described in Paper I to explain observations of the magnetic massive star Sco. The focus of this study is on the magnetic field amplification process, the magnetic field geometry, the mixing during the merger and the rotational configuration of the resulting merger product. The main findings can be summarised as follows.
@@
We observe robust magnetic-field amplification by more than ten orders of magnitude. The magnetic-field growth is initially sourced by turbulent energy in the merger and then by large-scale, ordered azimuthal fluid motions. In the end, the magnetic field energy reaches a super-equipartition level with respect to the kinetic energy in the radial and -direction, i.e. our proxy of the turbulent energy. The final outcome appears robust and does not depend on the initial separation of the binary, the numerical resolution and the strength of the seed magnetic field. We thus conclude that the memory of the initial seed magnetic field is lost.
The magnetic field is initially amplified by small-scale turbulent dynamos invoking the Kelvin-Helmholtz (shear) and magneto-rotational instabilities. Later, the large-scale, ordered azimuthal fluid motions drive another dynamo process that pumps the magnetic energy from small () to large () scales. The magnetic energy power spectrum in the final merger structure follows that of the kinetic energy and the merger remnant is threaded by a large-scale magnetic field.
The final magnetic-field geometry is characterised by a large-scale poloidal and toroidal structure with a still non-negligible small-scale component. The field structure resembles those found by Braithwaite & Spruit (2004) and Braithwaite & Nordlund (2006b) to be stable over thermal up to possibly nuclear timescales.
@@
This work is only the beginning of a detailed exploration of the complex physics in stellar mergers and many questions remain unanswered. The stability of the amplified magnetic fields needs to be assessed and the further evolution of the star–torus structure will determine the exact properties of merged stars such as their rotational velocities and surface chemical chemistry (see also Schneider et al. 2020). Addressing these questions is essential to further understand the properties of magnetic massive stars and their possible highly-magnetic white-dwarf and neutron-star remnants (cf. polars and magnetars).
@@ Acknowledgements.
- Ryu et al. (2024) Ryu, T., Sills, A., Pakmor, R., de Mink, S., & Mathieu, R. 2024, arXiv e-prints, arXiv:2410.00148

### v2

@@ 3.5 Dynamical ejecta
Because the further evolution of the merger product and also its optical display might be influenced by dynamical ejecta, we compute how much mass is unbound at the end of the simulation. To this end, we sum the mass of all cells in the simulation for which the total energy is positive. To get an upper limit, contributions from the potential, kinetic and also internal energies are taken into account. After the merger settled into the core–disk structure, at d, about of material is unbound, corresponding to 0.14% of the total mass of the system. Not much mass is ejected compared to other stellar mergers (e.g., common envelope phases, see Ohlmann et al. 2016a), also because the binding energy of the system is still quite large, with about . The unbound mass moves outwards with an average radial velocity of about 180 km/s and is predominantly ejected near the orbital plane with a distribution that shows some tails to the polar directions.
@@ 4.1 Impact of numerical parameters
The main numerical parameters that affect the simulation outcome are the resolution, the initial distance and the magnetic field configuration. To assess the impact of the resolution, we ran Model 1 with about 10 times more cells than Model 2 (cf. Table 2). Because the adaptive mesh leads to cells with nearly constant mass, this leads to a higher resolution in volume by a factor of 10, which roughly corresponds to a factor of 2 in each spatial dimension. Overall, these two simulations show quite similar behavior in their orbital evolution (Fig. 4). The evolution of the magnetic field amplification differs slightly for the two models (Fig. 6): the amplification sets in later for Model 2 with fewer cells. The final field strength and configuration, however, are quite similar, which demonstrates that they are robust with respect to the resolution.
The initial distance influences the outcome as more angular momentum is present in the system for larger initial distances. In principle, the simulation should start at Roche-lobe overflow; however, this is prohibitively expensive in computational time because the mass transfer progresses only slowly in the beginning, leading to a long time to the actual merger. As already described, we remove angular momentum artificially to obtain a starting configuration that leads to a merger more quickly. In reality, the system would have more angular momentum, leading probably to a longer mass loss via L2 and L3 in the phase before the actual merger. The largest impact of this is probably on the detailed structure of the outer part of the disk. To assess the impact of a larger initial separation, we ran Model 3 with a larger initial distance. As one can see in Fig. 4, it takes longer until the cores finally merge, but the overall evolution is similar. Also, the magnetic field evolution (Fig. 6) shows a slightly different shape, but the final field strength is of similar magnitude and configuration. Thus, we expect that the main conclusions are not affected by the detailed choice of this parameter, but that it mainly affects the outermost structure of the disk.
@@
A typical numerical issue in MHD simulations is fulfilling the divergence constraint . In principle, the continuous equations of MHD preserve this constraint for initial conditions fulfilling it. However, discretization errors also lead to errors in . As explained in Pakmor & Springel (2013), we use the Powell method (Powell et al. 1999) to control the errors in the divergence. We have checked that the errors in our simulations behave very similarly to those in the tests shown there (see Fig. 12): averages over larger regions are small and typically adjacent cells show different signs in the error, indicating that the large-scale impact of these local errors is small. The volume-weighted average of the relative divergence error is , , , and for the times given in Fig. 12.
@@ 4.2 Comparison to other simulations
Glebbeek et al. (2013) have conducted a series of stellar merger simulations as head-on collisions and subsequent stellar evolution calculations for a range of masses and evolutionary points. They used an SPH code with particles per star, whereas we use a grid-based approach with 1 to 2 million cells per star and also include MHD. By simulating head-on collisions, they avoid dealing with rotation in the merger product and also with angular momentum transport. The collision may lead to more heating compared to the merger in our simulations. Moreover, there is probably no disk forming around the merger product in their simulations due to the missing angular momentum. In their simulations, the equal-mass collisions eject between 4% and 8% of the total mass (see their Figure 3), whereas in our simulation, only 0.14% of the mass is unbound (compare Sect. 3.5). This might be due to the difference in the initial conditions as material can escape more easily in head-on collisions, whereas in our merger simulations a large disk forms that is bound. In our simulation, the core is dominated by material from the secondary star, corresponding to “case S” in Glebbeek et al. (2013). As explained there, this can be attributed to the lower buoyancy in the secondary star. They find “case S” for relatively unevolved primaries at moderate mass ratios, which would also apply to our case. For further differences between binary mergers and head-on collisions, see e.g. Schneider (2025) and Vynatheya et al. (2025).
@@
Glebbeek & Pols (2008) have analyzed stellar collision products. However, they do not conduct hydrodynamic simulations, but they use the entropy sorting method from Lombardi et al. (2002) to compute the merger products and then they focus on the further evolution. Lombardi et al. (2002) use results from earlier collision simulations using an SPH code to construct a model for predicting the outcome of such a collision. This model mainly sorts the material by its entropy, but takes into account more effects such as mass loss and shock heating. For our simulations, comparing the merger product with an entropy-sorted mix of the initial stellar models, one can clearly see similar effects: the entropy in the simulated merger is larger especially in the center and in the outermost parts. Simple entropy sorting of the initial stellar models does not provide the right structure of the merged stars. Shock heating, chemical mixing and changes in the total mass of the merged star and thereby in density and temperature modify the entropy of the original fluid elements. Still, the final merged star has an outward increasing entropy profile as is expected for buoyantly stable stratifications (Landau & Lifshitz 1987; Bisnovatyi-Kogan 2001). The models from Lombardi et al. (2002) are based on earlier collision simulations, amongst others the off-axis collisions conducted by Sills et al. (2001). They use an SPH code with about particles for the full simulation and target low-mass stars, mainly to study blue stragglers similar to Lombardi et al. (2002) and Glebbeek & Pols (2008). For further comparisons between simplified merger methods and the outcome of binary mergers and head-on collisions, see Heller et al. (2025). In particular, these authors use the outcome of Model 1 presented here in their comparison.
@@
Ryu et al. (2025) simulate parabolic head-on collisions of low-mass stars with non-zero impact parameters similar to those mentioned above, but they also use the moving-mesh MHD code Arepo and find strong magnetic field amplification. Such collisions contain less angular momentum than mergers in circular binaries studied here and thus form not so clear massive disk-like structures in the merger remnant as found here. The magnetic field amplification mechanism is otherwise in qualitative agreement with that found in our mergers of circular high-mass binaries: the head-on collisions invoke several contact/grazing encounters before the final coalescence, and, in each encounter, small-scale dynamos amplify the magnetic fields. From the work presented in Ryu et al. (2025), however, it remains unclear whether large-scale magnetic fields are formed. Vynatheya et al. (2025) compare head-on collisions of massive stars with binary mergers of the same stars. They find systematic differences between those; in particular, they find that merger remnants have signs of an ordering of the magnetic field similar to our simulation, which is absent for collision remnants. They focus on a qualitative analysis of the final state of their collision and merger remnants. The latter seem consistent with our simulation.
@@
Remarkably, the magnetic field amplification process in mergers of MS stars is so similar to that in mergers of white dwarfs and neutron stars presented in Pakmor et al. (2024) and Kiuchi et al. (2024), respectively. Also in these binary white-dwarf and neutron-star mergers, small-scale magnetic fields are produced by a turbulent dynamo and magnetic energy is then transferred to larger scales by a large-scale, MRI-driven dynamo. A similar situation is also found in neutron star–white dwarf mergers (Morán-Fraile et al. 2024) and common-envelope phases of low-mass and high-mass stars (Ondratschek et al. 2022; Vetter et al. 2024). Despite the similarities, all these simulations feature outflows that are magnetically driven, whereas our simulations do not show such outflows. This might be because we did not follow our simulations for long enough after the merger.
@@ 5 Conclusions
We have studied in detail the magnetic-field amplification in the merger of a and main sequence star in an isolated binary system and the properties of the resulting merged star. The simulations were first described in Paper I to explain observations of the magnetic massive star Sco. The focus of this study is on the magnetic field amplification process, the magnetic field geometry, the mixing during the merger and the rotational configuration of the resulting merger product. The main findings can be summarised as follows.
@@
We observe robust magnetic-field amplification by more than ten orders of magnitude. The magnetic-field growth is initially sourced by turbulent energy in the merger and then by large-scale, ordered azimuthal fluid motions. In the end, the magnetic field energy reaches a super-equipartition level with respect to the kinetic energy in the radial and -direction, i.e. our proxy of the turbulent energy. The outcome appears robust and does not depend on the initial separation of the binary, the numerical resolution and the strength of the seed magnetic field. We thus conclude that the memory of the initial seed magnetic field is lost.
The magnetic field is initially amplified by small-scale turbulent dynamos invoking the Kelvin-Helmholtz (shear) and magneto-rotational instabilities. Later, the large-scale, ordered azimuthal fluid motions drive another dynamo process that pumps the magnetic energy from small () to large () scales. The magnetic energy power spectrum in the final merger structure follows that of the kinetic energy, and the merger remnant is threaded by a large-scale magnetic field.
The final magnetic-field geometry is characterised by a large-scale poloidal and toroidal structure with a still non-negligible small-scale component. The field structure resembles that found by Braithwaite & Spruit (2004) and Braithwaite & Nordlund (2006b) to be stable over thermal up to possibly nuclear timescales.
@@
This work is only the beginning of a detailed exploration of the complex physics in stellar mergers, and several questions remain unanswered. The stability of the amplified magnetic fields needs to be assessed, and the further evolution of the star–torus structure will determine the exact properties of merged stars such as their rotational velocities and surface chemistry (see also Schneider et al. 2020; Schneider 2025). Addressing these questions is essential to further understand the properties of magnetic massive stars and their possible highly-magnetic white-dwarf and neutron-star remnants (cf. polars and magnetars).
@@ Acknowledgements.
- Bisnovatyi-Kogan (2001) Bisnovatyi-Kogan, G. S. 2001, Stellar physics. Vol.1: Fundamental concepts and stellar equilibrium
@@
- Heller et al. (2025) Heller, M., Schneider, F. R. N., Henneco, J., Bronner, V. A., & Lau, M. Y. M. 2025, arXiv e-prints, arXiv:2512.14669
@@
- Ryu et al. (2025) Ryu, T., Sills, A., Pakmor, R., de Mink, S., & Mathieu, R. 2025, ApJ, 980, L38